OpenStudy (anonymous):

A bullet of mass 50 g is moving with velocity 500 m/s . It penetrates 10m into a still target and comes to rest . Calculate (A) the KE possessed by the bullet (B) the avg. retarding force offered by the target .

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

\[K.E = \frac{ 1 }{ 2 } m * v^2\]

OpenStudy (anonymous):

and (B) ?

OpenStudy (anonymous):

Hint: Convert Mass gram to Kg

OpenStudy (anonymous):

For b u have to Find the Acceleration : \[v^2 - u^2 = 2as\]

OpenStudy (anonymous):

V = 0 Since the Bullet Comes to Rest

OpenStudy (anonymous):

and U is 500 ?

OpenStudy (anonymous):

\[K.E = \frac{ 1 }{2 } * 50 * 10^{-3} * 500 * 500\]

OpenStudy (anonymous):

\[V^2 - u^2 = 2as\] \[a = \frac{ u^2 }{ 2s }\]

OpenStudy (anonymous):

\[a= \frac{ 500 * 500 }{ 2 * 10 }\]

OpenStudy (anonymous):

After Finding a \[F = M*a\]

OpenStudy (anonymous):

hope...this helps u..)

OpenStudy (anonymous):

what is the meaning of avg. retarding force ?

OpenStudy (anonymous):

just force ?

OpenStudy (anonymous):

@yahoo there ?

OpenStudy (anonymous):

It is Just Force..)

OpenStudy (anonymous):

okay , ty :)

OpenStudy (anonymous):

Welcome...)